OpenStudy (anonymous):

Will receive medal for exact answer! The length of a diagonal of a large square garden is 15 yards. What is the approximate length of one side of the square garden? A. 8.6 yards B. 10.6 yards C. 21.2 yards D. 25.9 yards

OpenStudy (anonymous):



OpenStudy (anonymous):

Explanation and answer preferred please

OpenStudy (anonymous):

You could use pythagorean theorem: \[a ^{2}+b^{2}=c ^{2}\] c = 15 so work backwards

OpenStudy (anonymous):

let x be the edge of cube.Then x^2+x^2=15^2 2x^2=15^2 x^2=(15^2)/2=2(15/2)^2 x=(15/2)sqrt2=(15/2)(1.414)=15*0.707=10.605 yards